# SAEs Are Good for Steering – If You Select the Right Features

### v1

@@ 4.4 Evaluation on AxBench
AxBench was recently proposed as a dataset to evaluate steering methods Wu et al. (2025). They compare steering with SAE features to other methods (including supervised methods), and find SAE features relatively ineffective. However, we believe this is partially due to a non-principled selection of SAE features; we propose to remedy this using the output score.
We evaluate our findings on instruction-tuned Gemma2-9B using the Concept500 dataset of Wu et al. (2025), which includes 1000 <concept, SAE feature>-pairs from layers 20 and 31 of the model. We evaluate the steered texts using the metrics defined by Wu et al.: (1) the concept score measures if the concept was incorporated in the generated text; (2) the fluency score measures the coherency of the text; and (3) the instruct score measures the alignment of the generated text with the given instructions. For each feature, we compute the harmonic mean of the three metrics. See Appendix D for additional details on the steering setup, metrics, and baseline methods.
(a) Gemma2-2B.
(b) Gemma2-9B.
Figure 6: Relationship between input and output scores for features in Gemma2-2B and Gemma2-9B. Most features lie near the axes, indicating that most features are either input or output features—though a few are both.
Figure 7: Generation results when steering with features that have both high input and high output scores in Gemma2-2B. The top logit lens tokens (left; top-1 in blue) do not appear directly in the generated text (right). The steered tokens (magenta), appearing lower in the logit lens ranking, often have strong collocational associations with the top logit lens tokens.
We report the mean score on layers 20 and 31 of instruction-tuned Gemma2-9B (Table 1). We find a nearly threefold improvement in SAE steering scores relative to Wu et al.. Note that our replication of their experimental setting (without filtering) yields higher scores for SAEs; this can be a result of different sampled instructions per feature, [^2] Wu et al. (2025) sample 10 instructions per feature from pre-existing datasets, but do not release these instructions. We sample 10 instructions from the same datasets, which may be different compared to the sample of Wu et al.. or possibly due to evaluation instability introduced by the use of an external LLM. With output score filtering, SAE features top steering performance among the representation-based methods at L31; for L20, they get 90.7% of the performance of the best method. This is in contrast with the results of Wu et al., where SAEs significantly underperforms ReFT-r1, a weakly supervised method they propose as a competitive alternative to prompting. These results demonstrates that with effective feature selection SAEs are comparable with existing methods, including supervised or weakly-supervised methods which require concept-specific datasets.
@@ 5 The Relationship Between Input and Output Scores
We next examine how input and output scores interact. Figure 3 suggests that high input and output scores are rarely observed in the same layers, but do they sometimes co-occur in the same features?
Figure 6 visualizes the relationship between the two scores: Indeed, most samples cluster near the axes, exhibiting either a high output score with a near-zero input score, or vice-versa. However, there do exist hybrid features: features with both high input and high output scores.
Figure 7 illustrates generation results when steering with hybrid features from Gemma2-2B. Generated tokens that also appear in the feature’s top-20 logit lens tokens are highlighted in magenta. These are often not the top-ranked tokens under the logit lens, but they tend to rank moderately high and collocate with the top tokens (highlighted in blue). For instance, for feature 6820 in layer 18, “contact” is the top logit lens token, but the output text repeatedly includes “lenses”, a token that appears further down the list but that is semantically and syntactically related.
(a) Gemma2-2B.
(b) Gemma2-9B.
@@
We verify this intuition by quantifying collocation patterns between generated tokens and top logit lens tokens using their pointwise mutual information (PMI) [^3] We use the pre-computed PMI over the webtext corpus in NLTK Bird and Loper (2004), and only include token pairs that have this pre-computed score (100–200 pairs per model).. A PMI of zero indicates that two tokens co-occur no more frequently than chance, and negative and positive scores indicate lower- or higher-than-chance co-occurrence, respectively.
Features with high output score ($S_{out}\geq 0.1$) and low input score ($S_{in}<0.1$) have a negative PMI value on average (Figure 8). This can be attributed to the high generation success of these features, i.e., that the top logit lens token is equal to the generated token. In most cases, two instances of the same token are not likely to consecutively co-occur, thus obtaining a low PMI score.
@@ 6.1 Stages of Processing in LMs
The different stages of processing within NLP models have long been studied Belinkov et al. (2017); Zhang and Bowman (2018); Liu et al. (2019); Brunner et al. (2020). In transformer-based LMs, a large body of work shows that different properties emerge in different layers. Early layers focus on syntactic tasks such as POS, while semantic information appears in later layers Tenney et al. (2019); Elazar et al. (2021); Geva et al. (2021). More recent work has demonstrated that intermediate layers are responsible for retrieving factual knowledge and enriching latent representations Meng et al. (2022); Geva et al. (2023); Hernandez et al. (2024).
Another line of work has focused on so-called prediction neurons, which increase the probability of coherent sets of tokens. This work characterizes prediction neurons by properties of their logit lens distribution Gurnee et al. (2024); Lad et al. (2024); Bloom and Lin (2024). In contrast, our output score directly measures the causal effect of a feature on predicting a pre-defined set of tokens via counterfactual interventions. Relatedly, Lad et al. (2024) have shown that neurons in early layers pay more attention to input tokens in their proximity compared to later layers, while prediction neurons emerge later, after about 50% of the model depth, in line with our findings on SAEs.
A closely related line of work aims to explain SAE features in natural language—for instance, by feeding inputs and activations into an external LLM Bills et al. (2023); Huben et al. (2024). However, this method results in errors in both precision and recall (Huang et al., 2023), and negatively (albeit weakly) correlates with their causal role on average (Paulo et al., 2024). Gur-Arieh et al. (2025) suggest that SAE features are better explained in terms of their activations and their effect on the output(as quantified by projecting features into vocabulary space). In this work, our aim is not to explain features, but rather categorize them with respect to their usefulness for steering. Additionally, we differentiate between two key feature roles (often mutually exclusive); this helps explain these prior findings and failure cases.
@@ 6.2 Steering LMs
Many approaches exist for precisely influencing the outputs generated by LMs Zou et al. (2023). These include prompt engineering Wu et al. (2025); Taveekitworachai et al. (2024), steering vectors Subramani et al. (2022); Teehan et al. (2022); Liu et al. (2023) or inference-time interventions on activations Turner et al. (2023); van der Weij et al. (2024); Rimsky et al. (2024). Steering was shown to be useful not only for directing generated content to a specific topic or concept, but also for style transfer Lai et al. (2024), mitigating hallucinations Li et al. (2023a); Simhi et al. (2024), and debiasing (Li et al., 2025).
While most work in this area steers via interventions on full hidden states, earlier work attempted to influence model behavior by intervening on small sets of neurons Bau et al. (2019). However, the polysemanticity of neurons makes them poor candidates for effective steering Bricken et al. (2023). In contrast, SAEs were shown to result in meaningful steering towards human-understandable concepts; a famous example involved steering toward responses related to the Golden Gate Bridge Templeton et al. (2024), and another involved amplifying or mitigating social and political biases Durmus et al. (2024); Marks et al. (2025). Recently, Wu et al. (2025) evaluated SAE steering against many methods, including supervised methods such as full fine-tuning, prompting, and difference-in-means Larsen et al. (2016). They found that even these simple baselines outperform SAEs. However, our work shows that most of the gap can be closed via more careful choice of SAE features. Typical work chooses features for steering based on natural language explanations generated based on each feature’s activation patterns Huben et al. (2024); Durmus et al. (2024); our findings instead suggest that influence on output is a better proxy for steering efficacy, and that input activations have little predictive power for finding good steering features.
@@ 7 Conclusions
Following previous work, we focus our analysis on the Gemma2 family of models and the Gemma-Scope SAEs Team et al. (2024); Lieberum et al. (2024). While these models are widely adopted for interpretability research, further analysis across different LMs and SAE architectures would be beneficial to assess the robustness of our method.
Third, our method focuses on steering using a single SAE feature. In practice, interactions between features may lead to better and more complex effects on generation. Understanding how multiple features combine or interfere remains an open challenge.
@@ Acknowledgments
This research was supported by an Azrieli Foundation Early Career Faculty Fellowship and Open Philanthropy. Dana Arad is supported by the Ariane de Rothschild Women Doctoral Program. Aaron Mueller is supported by a postdoctoral fellowship under the Zuckerman STEM Leadership Program. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Additional Steering Examples
Table 2 shows additional examples of steering. Features with high output scores result in meaningful steering, while features with high input score and low output score do not have any visible effect on the generated text.
@@
Table 2: Examples of steering with features with different output and input score values.
Appendix B Identifying Input Features
Table 3 shows examples of features from layers 0 and 1 having high input scores. For each feature, the table includes it’s top-5 logit lens tokens as well as examples for input texts that activated this feature. The tokens where the feature activated most strongly are marked using an underline.
@@ Appendix B Identifying Input Features
Appendix C Steering Details
For our main experiments, we test steering factor values $s\in\{0.2,0.4,0.8,1.2,1.6,2.0,3.0,4.0,\\ 6.0,8.0,10.0,20.0\}$. We generate 20 tokens using a temperature of $0.7$ after each of the prefixes listed in Table 4.
@@ Appendix C Steering Details
Appendix D AxBench Details
@@
Each concept in the Concept500 dataset is annotated as either ”text”, ”math”, or ”code”. Following their setup we randomly sample 10 instructions per concept from instruction datasets that match the concept genre: Free Dolly dataset for text instructions Conover et al. (2023), GSM8K for math Cobbe et al. (2021), and Alpaca-Eval for code Li et al. (2023b).
@@
We evaluate the steered texts using the metrics defined by Wu et al.: (1) the concept score ($cs$) measures if the concept was incorporated in the generated text, (2) the fluency score ($fs$) measures the coherency of the text, and (3) the instruct score ($is$) measures the alignment of the generated text with the given instructions. For each metric ($m\in\{\text{cs},\text{fs},\text{is}\}$) and steered text $s$, an external rater returns a discrete score of either 0, 1, or 2: $m(s)\in\{0,1,2\}$. As the external LLM rater, we use Claude 3.7 sonnet (2025-02-19) Anthropic (2024). The prompts for all metrics are given in Tables 5, 6, and7.
@@
Table 1 shows the reported results of steering using various methods, as achieved by Wu et al. (2025). We provide additional details on each of these methods:
@@
  • LoRA. A parameter-efficient finetuning method Xu et al. (2024), trained to minimize the language modeling loss on a set of positive examples that contain the concept.
  • LoReFT. Low-rank representation finetuning Wu et al. (2024), with the same objective as LoRA.
  • ReFT-r1. A weakly supervised approach based on Wu et al. (2024) Wu et al. (2025). ReFT-r1 jointly learns concept detection and steering on supervised data by combining the training objectives of linear probing and supervised steering.
  • Difference-in-means. Difference-in-means (DiffMean) Larsen et al. (2016) uses the difference between averaged representations from two classes of inputs as a steering vector.
  • SAE-A. SAEs with AUROC selection Wu et al. (2025). For each feature, its max-pooled activations are computed over a set of positive training examples. Top-scoring features, based on the AUROC over the dataset given true labels, are used for steering.
@@
Appendix E Computational Requirements
The experiments in this work were conducted using an NVIDIA A40404040 node with $8$ $48$GB GPUs. The complete set of experiments, including initial exploration, consumed 300-500 GPU hours.
Computing the output score for a single feature takes 6 seconds on average on a single A40404040 $48$GB GPU. Computing the input scores given pre-computed activations takes negligible time.

### v2

@@ 4.4 Evaluation on AxBench
AxBench was recently proposed as a dataset to evaluate steering methods wu2025axbench. They compare steering with SAE features to other methods (including supervised methods), and find SAE features relatively ineffective. However, we believe this is partially due to a non-principled selection of SAE features; we propose to remedy this using the output score.
[…]
(a) Gemma-2-2B.
(b) Gemma-2-9B.
Figure 6: Relationship between input and output scores for features in Gemma-2-2B and Gemma-2-9B. Most features lie near the axes, indicating that most features are either input or output features—though a few are both.
Figure 7: Generation results when steering with features that have both high input and high output scores in Gemma-2-2B. The top logit lens tokens (left; top-1 in blue) do not appear directly in the generated text (right). The steered tokens (magenta), appearing lower in the logit lens ranking, often have strong collocational associations with the top logit lens tokens.
We report the mean score on layers 20 and 31 of instruction-tuned Gemma-2-9B (Table˜1). We find a nearly threefold improvement in SAE steering scores relative to wu2025axbench. Note that our replication of their experimental setting (without filtering) yields higher scores for SAEs; this can be a result of different sampled instructions per feature,222wu2025axbench sample 10 instructions per feature from pre-existing datasets, but do not release these instructions. We sample 10 instructions from the same datasets, which may be different compared to the sample of wu2025axbench. or possibly due to evaluation instability introduced by the use of an external LLM. With output score filtering, SAE features top steering performance among the representation-based methods at L31; for L20, they get 90.7% of the performance of the best method. This is in contrast with the results of wu2025axbench, where SAEs significantly underperforms ReFT-r1, a weakly supervised method they propose as a competitive alternative to prompting. These results demonstrates that with effective feature selection SAEs are comparable with existing methods, including supervised or weakly-supervised methods which require concept-specific datasets.
@@ 5 The Relationship Between Input and Output Scores
We next examine how input and output scores interact. Figure˜3 suggests that high input and output scores are rarely observed in the same layers, but do they sometimes co-occur in the same features?
Figure˜6 visualizes the relationship between the two scores: Indeed, most samples cluster near the axes, exhibiting either a high output score with a near-zero input score, or vice-versa. However, there do exist hybrid features: features with both high input and high output scores.
Figure˜7 illustrates generation results when steering with hybrid features from Gemma-2-2B. Generated tokens that also appear in the feature’s top-20 logit lens tokens are highlighted in magenta. These are often not the top-ranked tokens under the logit lens, but they tend to rank moderately high and collocate with the top tokens (highlighted in blue). For instance, for feature 6820 in layer 18, “contact” is the top logit lens token, but the output text repeatedly includes “lenses”, a token that appears further down the list but that is semantically and syntactically related.
(a) Gemma-2-2B.
(b) Gemma-2-9B.
@@
We verify this intuition by quantifying collocation patterns between generated tokens and top logit lens tokens using their pointwise mutual information (PMI) [^3] We use the pre-computed PMI over the webtext corpus in NLTK bird-loper-2004-nltk, and only include token pairs that have this pre-computed score (100–200 pairs per model).. A PMI of zero indicates that two tokens co-occur no more frequently than chance, and negative and positive scores indicate lower- or higher-than-chance co-occurrence, respectively.
Features with high output score ($S_{out}\geq 0.1$) and low input score ($S_{in}<0.1$) have a negative PMI value on average (Figure˜8). This can be attributed to the high generation success of these features, i.e., that the top logit lens token is equal to the generated token. In most cases, two instances of the same token are not likely to consecutively co-occur, thus obtaining a low PMI score.
@@ 6.1 Stages of Processing in LMs
The different stages of processing within NLP models have long been studied belinkov2017translation; zhang2018language; liu2019linguistic; brunner2020identifiability. In transformer-based LMs, a large body of work shows that different properties emerge in different layers. Early layers focus on syntactic tasks such as POS, while semantic information appears in later layers bert2019tenney; elazar2021amnesic; geva2021keyvalue. More recent work has demonstrated that intermediate layers are responsible for retrieving factual knowledge and enriching latent representations meng2022locating; geva2023recall; hernandez2024linearity; arad2024refact.
Another line of work has focused on so-called prediction neurons, which increase the probability of coherent sets of tokens. This work characterizes prediction neurons by properties of their logit lens distribution gurnee2024universal; lad2024mstages; saelogitlens. In contrast, our output score directly measures the causal effect of a feature on predicting a pre-defined set of tokens via counterfactual interventions. Relatedly, lad2024mstages have shown that neurons in early layers pay more attention to input tokens in their proximity compared to later layers, while prediction neurons emerge later, after about 50% of the model depth, in line with our findings on SAEs.
A closely related line of work aims to explain SAE features in natural language—for instance, by feeding inputs and activations into an external LLM bills2024explaining; interpretable2024huben. However, this method results in errors in both precision and recall (huang-etal-2023-rigorously), and negatively (albeit weakly) correlates with their causal role on average (paulo2024automatically). gurarieh2025enhancing suggest that SAE features are better explained in terms of their activations and their effect on the output(as quantified by projecting features into vocabulary space). In this work, our aim is not to explain features, but rather categorize them with respect to their usefulness for steering. Additionally, we differentiate between two key feature roles (often mutually exclusive); this helps explain these prior findings and failure cases.
@@ 6.2 Steering LMs
Many approaches exist for precisely influencing the outputs generated by LMs zou2023representation. These include prompt engineering wu2025axbench; taveekitworachai2024null, steering vectors subramani2022extracting; teehan2022emergent; liu2023context or inference-time interventions on activations turner2023steering; van2024extending; rimsky2024steering. Steering was shown to be useful not only for directing generated content to a specific topic or concept, but also for style transfer lai2024style, mitigating hallucinations li2023inference; simhi2024constructing, and debiasing (li2025fairsteer).
While most work in this area steers via interventions on full hidden states, earlier work attempted to influence model behavior by intervening on small sets of neurons bau2018identifying. However, the polysemanticity of neurons makes them poor candidates for effective steering bricken2023monosemanticity. In contrast, SAEs were shown to result in meaningful steering towards human-understandable concepts; a famous example involved steering toward responses related to the Golden Gate Bridge scaling2024templeton, and another involved amplifying or mitigating social and political biases durmus2024steering; marks2025sparse. Recently, wu2025axbench evaluated SAE steering against many methods, including supervised methods such as full fine-tuning, prompting, and difference-in-means pmlr-v48-larsen16. They found that even these simple baselines outperform SAEs. However, our work shows that most of the gap can be closed via more careful choice of SAE features. Typical work chooses features for steering based on natural language explanations generated based on each feature’s activation patterns interpretable2024huben; durmus2024steering; our findings instead suggest that influence on output is a better proxy for steering efficacy, and that input activations have little predictive power for finding good steering features.
@@ 7 Conclusions
Additionally, our method focuses on steering using a single SAE feature. In practice, interactions between features may lead to better and more complex effects on generation (wattenberg2024relational; singhvi-etal-2025-using). Understanding how multiple features combine or interfere remains an open challenge.
@@ Acknowledgments
This research was supported by an Azrieli Foundation Early Career Faculty Fellowship and by Open Philanthropy. Dana Arad is supported by the Ariane de Rothschild Women Doctoral Program. Aaron Mueller was supported by a postdoctoral fellowship under the Zuckerman STEM Leadership Program. This research was funded by the European Union (ERC, Control-LM, 101165402). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.
@@ Appendix A Additional Steering Examples
Tables 2, 3, and 4 show examples of steering Gemma-2-2B, Pythia, and Llama-3.1, respectively. Features with high output scores result in meaningful steering, while features with high input score and low output score do not have any visible effect on the generated text.
@@
Table 2: Examples of steering with features with different output and input score values in Gemma-2-2B.
Layer Feature
Input
Score
Output
Score
Top-5 Logit Lens Tokens
Optimal
Steering
Factor
Generated Text
4 7772 0 0.794
[’_Firefox’, ’_Chrome’,
’_browser’, ’_Mozilla’,
’_browsers’]
1.2
Findings show that a
site may play a key role
in the development of a
web browser.
5 21568 0 0.474
[’_Barack’, ’_Obama’,
’_Donald’, ’_Trump’,
’_Bush’]
0.8
The legend goes that the guy
was a great- Barack Obama,
he’d probably be the biggest
supporter of Barack Obama ever
Table 3: Examples of steering with features with different output and input score values in Pythia-70m.
Layer Feature
Input
Score
Output
Score
Top-5 Logit Lens Tokens
Optimal
Steering
Factor
Generated Text
2 25580 0.436 0
[’_sources’, ’_SOUR’,
’sources’, ’iped’,
’/source’]
0.8
That reminds me of the time
when I was in a public of knowledge
for the answer of information that said
the information was said that said
20 17816 0 0.669
[’_visa’, ’_immigration’,
’_Immigration’, ’_visas’,
’_Visa’]
1.6
I believe that the H-1B visa program
is an important tool for employers
to access the best talent to fill
26 21627 0 0.665
[’_disability’, ’_disabled’,
’_Disability’, ’_disable’,
’_Disabled’]
1.2
Findings show that the mainstream
media has a strong impact on public
opinion on disability.
Table 4: Examples of steering with features with different output and input score values in Llama-3.1-8B.
Appendix B Results on Llama-3.1-8B and Pythia-70m
Figure˜9 demonstrates the distributions of scores across layers for Llama-3.1 and Pythia.
Figure˜10 shows the mean generation success @ 20 of steered generations when filtering out features with output scores below varying thresholds (magenta) on Llama-3.1 and Pythia. Similarly to the Gemma models, as the threshold increases, performance improves steadily, indicating that features with higher output scores consistently lead to more successful steering.
(a) Llama-3.1-8B.
(b) Pythia-70m.
Figure 9: Input and output scores across layers in Llama-3.1-8B and Pythia-70m. The solid lines represent the median input score (blue) and output score (magenta), while the shaded regions denote the interquartile range (25th to 75th percentile), capturing the variability across features within each layer. In these models we observe that high output scores emerge in later layers, while input score is mostly zero across all layers.
(a) Llama-3.1-8B.
(b) Pythia-70m.
Appendix C Identifying Input Features
Table˜5 shows examples of features from layers 0 and 1 having high input scores. For each feature, the table includes it’s top-5 logit lens tokens as well as examples for input texts that activated this feature. The tokens where the feature activated most strongly are marked using an underline.
@@ Appendix C Identifying Input Features
Appendix D Neutral Prompt Selection
Our output score relies on the use of a single neutral prompt as a prefix to the model’s generation. To evaluate the robustness of our score to the specific choice or neutral prompt we randomly selected 10 features from each layer of Gemma-2-2B (240 features overall) and computed their output scores using all $50$ of the neutral prompts in Table˜6. We find that the correlation between the output score computed with the original prompt and the mean score is $0.9557$, which is extremely high and indicates that the exact choice of the prompt has almost no impact on the results.
Appendix E Steering Details
For our main experiments, we test steering factor values $s\in\{0.2,0.4,0.8,1.2,1.6,2.0,3.0,4.0,\\ 6.0,8.0,10.0,20.0\}$. We generate 20 tokens using a temperature of $0.7$ after each of the prefixes listed in Table˜6.
@@ Appendix E Steering Details
Appendix F AxBench Details
@@
Each concept in the Concept500 dataset is annotated as either ”text”, ”math”, or ”code”. Following their setup we randomly sample 10 instructions per concept from instruction datasets that match the concept genre: Free Dolly dataset for text instructions DatabricksBlog2023DollyV2, GSM8K for math cobbe2021gsm8k, and Alpaca-Eval for code alpaca_eval.
@@
We evaluate the steered texts using the metrics defined by wu2025axbench: (1) the concept score ($cs$) measures if the concept was incorporated in the generated text, (2) the fluency score ($fs$) measures the coherency of the text, and (3) the instruct score ($is$) measures the alignment of the generated text with the given instructions. For each metric ($m\in\{\text{cs},\text{fs},\text{is}\}$) and steered text $s$, an external rater returns a discrete score of either 0, 1, or 2: $m(s)\in\{0,1,2\}$. As the external LLM rater, we use Claude 3.7 sonnet (2025-02-19) claude2024. The prompts for all metrics are given in Tables 7, 8, and9.
@@
Table˜1 shows the reported results of steering using various methods, as achieved by wu2025axbench. We provide additional details on each of these methods:
@@
  • LoRA. A parameter-efficient finetuning method xu2024lora, trained to minimize the language modeling loss on a set of positive examples that contain the concept.
  • LoReFT. Low-rank representation finetuning wu2024reft, with the same objective as LoRA.
  • ReFT-r1. A weakly supervised approach based on wu2024reft wu2025axbench. ReFT-r1 jointly learns concept detection and steering on supervised data by combining the training objectives of linear probing and supervised steering.
  • Difference-in-means. Difference-in-means (DiffMean) pmlr-v48-larsen16 uses the difference between averaged representations from two classes of inputs as a steering vector.
  • SAE-A. SAEs with AUROC selection wu2025axbench. For each feature, its max-pooled activations are computed over a set of positive training examples. Top-scoring features, based on the AUROC over the dataset given true labels, are used for steering.
@@
Appendix G Computational Requirements
The experiments in this work were conducted using an NVIDIA A4040 node with 88 4848GB GPUs. The complete set of experiments, including initial exploration, consumed 300-500 GPU hours.
Computing the output score for a single feature takes 6 seconds on average on a single A4040 4848GB GPU. Computing the input scores given pre-computed activations takes negligible time.